OpenStudy (anonymous):

How much water should be added to 1.00 mL of a 3.50 M solution of HCl to obtain a 2.00 M solution? A) 7 mL B) 1.75 mL C) .57 mL D) .75 mL E) Not enough information

OpenStudy (anonymous):

c=n/V, meaning that you have 3.5=n/0.001 and n=0.0035 mol, so the second solution will be 2=0.0035/V, meaning V=0.00175, which is 1.75 ml, so you'll have to put the difference between the two solutions 1.75-1.00=0.75 ml ==> D